OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 19 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

Ok to solve this one

OpenStudy (anonymous):

We can use the elimination method

OpenStudy (anonymous):

Ok. How do i do it though

OpenStudy (anonymous):

Ok here I will make sure you get it

OpenStudy (anonymous):

You see how there is -y and +y?

OpenStudy (anonymous):

Yes i do

OpenStudy (anonymous):

Ok that gives you the sign of being able to use elimination method

OpenStudy (anonymous):

So you just cross them, out or "Eliminate them"

OpenStudy (anonymous):

Then you add the x's

OpenStudy (anonymous):

you have 2x and x add them and get 3x

OpenStudy (anonymous):

Next since we already "eliminated the y's"

OpenStudy (anonymous):

We have = 19 = 2

OpenStudy (anonymous):

Just add those and get 21

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So you would then have 3x=21

OpenStudy (anonymous):

and you would then just solve it normally you would get x=7

OpenStudy (anonymous):

Which you get from dividing 21 by 3

OpenStudy (anonymous):

Now we have our x

OpenStudy (anonymous):

We still need to solve for our y

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

You can input our newly found x into any of the equations

OpenStudy (anonymous):

Which one you want to use? x - y = 19 2x + y = 2

OpenStudy (anonymous):

You can use any of them

OpenStudy (anonymous):

or the easier one

OpenStudy (anonymous):

So I pick the second one

OpenStudy (anonymous):

im not sure what you mean i just need to answer that Question....

OpenStudy (anonymous):

Ok so remember when we got our x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We still need to get our y

OpenStudy (anonymous):

So we have to replace x with our 7, you can do this with any of the equations in the questions

OpenStudy (anonymous):

I picked the second one because I like to multiply through

OpenStudy (anonymous):

2x + y = 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Is one I picked

OpenStudy (anonymous):

So 2*7= 14

OpenStudy (anonymous):

14 + y= 2

OpenStudy (anonymous):

Ok now subtract 14

OpenStudy (anonymous):

to 2

OpenStudy (anonymous):

and then your y will =-12

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

So your answer is (7,-12)

OpenStudy (anonymous):

Oh thank you !!!

OpenStudy (anonymous):

or if you want to write it out it is x=7 and y= -12

OpenStudy (anonymous):

Yeah no problem!

OpenStudy (anonymous):

Do you have any questions?

OpenStudy (anonymous):

Not at the moment ill get back to ya on that

OpenStudy (anonymous):

Sure, no problem!

OpenStudy (anonymous):

Auditions for a play in New York were held with 120 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many men auditioned? 44 40 76 80

OpenStudy (anonymous):

Ok this one

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2x-3(120-x)=20

OpenStudy (anonymous):

We made this from the question

OpenStudy (anonymous):

Now we just solve it

OpenStudy (anonymous):

So to get the rest we use distrubute property

OpenStudy (anonymous):

2x - 360 + 3x = 20

OpenStudy (anonymous):

So now we do + 360

OpenStudy (anonymous):

and combine 2x and 3x

OpenStudy (anonymous):

So you would get 5x= 380

OpenStudy (anonymous):

and just divide

OpenStudy (anonymous):

To get 76

OpenStudy (anonymous):

So 76 men auditioned

OpenStudy (anonymous):

Thank you again if i have another which i will ill type it in here

OpenStudy (anonymous):

Now we need to find the woman

OpenStudy (anonymous):

Wait we are not done yet

OpenStudy (anonymous):

So we just put in 76 for x

OpenStudy (anonymous):

120 - 76

OpenStudy (anonymous):

i only need to know how many men auditioned

OpenStudy (anonymous):

44 woman auditioned

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

Yeah no problem!

OpenStudy (anonymous):

The graph of a system of parallel lines has no solutions. Always Sometimes Never

OpenStudy (anonymous):

Never

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

It's Always

OpenStudy (anonymous):

It was one of those trick questions

OpenStudy (anonymous):

how sure because i know on a graph parallel lines have no soultion so i think it is never

OpenStudy (anonymous):

Always = Agree with the statement

OpenStudy (anonymous):

duh see i almost missed that question my bad.

OpenStudy (anonymous):

If you feel you are correct go with what you think

OpenStudy (anonymous):

no your right i didnt read it right

OpenStudy (anonymous):

Yeah I didnt read it right at first either

OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10 -2y + 10 2y + 10 2y - 10 -2y - 10

OpenStudy (anonymous):

this one

OpenStudy (anonymous):

Ok sorry about that I was solving it on a paper

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Ok so 3x= 8 -y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3x = 8 -y

OpenStudy (anonymous):

? ok so whats the answer?

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

@dpasingh Could you come here and help him, I am cooking right now

OpenStudy (anonymous):

Someone will come and help you give me a second

OpenStudy (anonymous):

@amistre64 Can you help him Im busy with other things right now

OpenStudy (amistre64):

they asked the same type of question that you did all the work for the first time; they keep on asking if "thats the answer" and "what is the answer" which tells me that they are not here to study any thing but to just have someone had them an answer to it .. so id really rather not :/ This is also why we would perfer that the asker participate in the solution process as opposed to being some by-stander and not having to do anything. We would also perfer that they close old questions that have been processed and post a new question instead of trying to fit them all into one post.

OpenStudy (anonymous):

Ok, thanks amistre